# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ 5.3. Fast kernel computations and efficient updates
    where ${\mathsf{V}}^{(m)}\overline{{\mathsf{V}}^{(m)}}={\mathsf{I}}$ and the first column of ${\mathsf{V}}^{(m)}$, denoted ${\boldsymbol{v}}_{1}^{(m)}$, is the constant $1/\sqrt{n}$
  2. (2)
    $V^{(m)}\in{\mathbb{C}}^{n\times n}$ satisfies ${\mathsf{V}}^{(0)}=(1)$ and
@@
    ${\mathsf{V}}^{(m+1)}=\begin{pmatrix}{\mathsf{V}}^{(m)}&{\mathsf{V}}^{(m)}\\ {\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}&-{\mathsf{V}}^{(m)}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}^{(m)})}\end{pmatrix}/\sqrt{2}.$
  3. (3)
    ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ are each computable at cost ${\mathcal{O}}(2^{m}m)$.
@@
For the case of lattice points with shift-invariant kernels and resulting circulant Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{T}}^{(m)}/\sqrt{2^{m}}$ is the scaled and permuted Fourier matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{\exp(-\pi\sqrt{-1}/2^{m}i)\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a bit-reversed FFT and bit-reversed IFFT respectively. For the case of digital nets with digitally-shift-invariant kernels and resulting nested block Toeplitz Gram matrices, ${\mathsf{V}}^{(m)}={\mathsf{H}}^{(m)}/\sqrt{n}$ is the scaled Hadamard matrix and ${\widetilde{\boldsymbol{w}}}^{(m)}=\{1\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}^{(m)}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{y}}$ can be computed using a FWHT and inverse FWHT respectively.
For ${\mathsf{\Lambda}}=\mathrm{diag}({\boldsymbol{\lambda}})$, we have
${\boldsymbol{\lambda}}=\sqrt{n}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}\left({\mathsf{V}}^{(m)}{\mathsf{\Lambda}}^{(m)}\overline{{\boldsymbol{v}}_{1}^{(m)}}\right)=\sqrt{n}\overline{{\mathsf{V}}^{(m)}}{\boldsymbol{k}}_{1}$
which can be computed at ${\mathcal{O}}(n\log n)$ cost. Moreover,
${\mathsf{K}}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}\odot{\boldsymbol{\lambda}})\qquad\text{and}\qquad{\mathsf{K}}^{-1}{\boldsymbol{y}}={\mathsf{V}}^{(m)}({\widetilde{\boldsymbol{y}}}./{\boldsymbol{\lambda}})$
@@
may each be evaluated at cost ${\mathcal{O}}(n\log n)$ where $\odot$ denotes the Hadamard (component-wise) product and ././. / denotes component-wise division.
To update the solutions of the motivating problems, it is easily shown that
${\mathsf{K}}^{(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\sqrt{2}\begin{pmatrix}{\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\\ {\mathsf{V}}^{(m)}\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}+{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)\end{pmatrix}$
${\mathsf{K}}^{-(m+1)}{\boldsymbol{y}}^{(m+1)}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}{\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\\ {\mathsf{V}}^{(m)}\left(\left({\widetilde{\boldsymbol{y}}}^{(2)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{y}}}^{(1)}\odot{\boldsymbol{\lambda}}^{(2)}\right)./{\widetilde{\boldsymbol{\gamma}}}\right)\end{pmatrix}$
where ${\widetilde{\boldsymbol{\gamma}}}={\boldsymbol{\lambda}}^{(1)}\odot{\boldsymbol{\lambda}}^{(1)}-{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\widetilde{\boldsymbol{w}}}^{(m)}\odot{\boldsymbol{\lambda}}^{(2)}\odot{\boldsymbol{\lambda}}^{(2)}$.
6. Numerical experiments
@@
All numerical experiments were carried out on a 2019 MacBook Pro with an Intel Core i7 processor. Figure 5 compares the wall-clock time required to generate point sets and perform fast transforms. Since scaling in the number dimensions and number of randomizations is linear, point sets are only generated in $d=1$ with a single randomization. IID points, lattices with shifts, and digital nets with LMS and digital shifts (including higher-order versions) are the fastest sequences to generate. Digital nets in base $b=2$ exploit Graycode order, integer storage of bit-vectors, and exclusive or (XOR) operations to perform digital addition. Generators of Halton point sets are slower to generate as they cannot exploit these advantages. NUS, especially higher order versions, are significantly slower to generate then LMS randomizations. Even so, higher-order LMS scrambling with digital shifts are sufficient to achieve higher-order RMSE convergence as we shown in the next experiment.
Figure 5. Comparison of time required to generate IID and low-discrepancy point sets and perform fast transforms. All point sets are generated in a single dimension and with a single randomizations. BR indicates the bit-reversed FFT and IFFT.
@@
Figure 6 shows higher-order LMS with digital shifts for base $b=2$ digital nets achieve higher-order RMSE convergence for RQMC methods. Specifically, we are able to achieve an RMSE of ${\mathcal{O}}(n^{-\min\{\alpha,{\widetilde{\alpha}}\}-1/2-\delta})$ where $\alpha$ is the higher-order digital interlacing of the net, ${\widetilde{\alpha}}$ is the smoothness of the integrand, and $\delta>0$ is arbitrarily small. The first two integrands reproduce the experiments of (Dick, 2011). The next two problems are test functions from Derek Bingham’s Virtual Library of Simulation Experiments (VLSE) [^14] https://www.sfu.ca/~ssurjano/uq.html. The last two Genz functions are used for forward uncertainty quantification in the Dakota software package (Adams et al., 2020) among other places. For each problem, the RMSE of the (Q)MC estimator (1) with ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$ is approximated using $300$ independent randomizations of a LD (or IID) point sets. For lattices, we periodize the integrand using a Baker-transform i.e. we use $\tilde{f}({\boldsymbol{x}})=f(1-2\lvert{\boldsymbol{x}}-1/2\rvert)$ in place of $f({\boldsymbol{x}})$ in (1). The Baker-transform does not change the expectation. A description of the integrands follows
Simple function, $d=1$:
has $f(X)=Xe^{X}-1$. This was used in (Dick, 2011) where higher-order digital net scrambling was first proposed.
Simple function, $d=2$:
has $f({\boldsymbol{X}})=X_{2}e^{X_{1}X_{2}}/(e-2)-1$. This was also used in (Dick, 2011).
Oakley & O’Hagan, $d=2$:
has $f({\boldsymbol{X}})=g(({\boldsymbol{X}}-1/2)/50)$ for $g({\boldsymbol{T}})=5+T_{1}+T_{2}+2\cos(T_{1})+2\cos(T_{2})$, see (Oakley and O’hagan, 2002) or the VLSE14.
G-Function, $d=3$:
has $f({\boldsymbol{X}})=\prod_{j=1}^{d}\frac{\lvert 4X_{j}-2\rvert-a_{j}}{1+a_{j}}$ with $a_{j}=(j-2)/2$ for $1\leq j\leq d$, see (Crestaux et al., 2007; Marrel et al., 2009) or the VLSE14.
Oscillatory Genz, $d=3$:
has $f({\boldsymbol{X}})=\cos\left(-\sum_{j=1}^{d}c_{j}X_{j}\right)$ with coefficients of the third kind $c_{j}=\exp\left(k\log\left(10^{-8}\right)/d\right)$. This is a common test function for uncertainty quantification which is is available in the Dakota software (Adams et al., 2020) among other.
Corner-peak Genz, $d=3$:
has $f({\boldsymbol{x}})=\left(1+\sum_{j=1}^{d}c_{j}X_{j}\right)^{-(d+1)}$ with coefficients of the second kind $c_{j}=1/j^{2}$. This is also used in the Dakota software (Adams et al., 2020) among others.
@@
Figure 6. The RMSE of the RQMC estimate for a few different integrands. Higher-order digital nets with linear matrix scrambling and digital shifts achieve higher-order convergence.
[…]
7. Conclusion and future work
@@
This work has reviewed routines for generating low-discrepancy point sets, randomizing them, and applying them to fast kernel methods. We have reviewed lattices, digital nets, and Halton point sets with randomizations spanning shifts, digital permutations, digital shifts, linear matrix scrambling, and nested uniform scrambling. Higher-order scramblings for digital nets were also considered. For kernel methods, we presented classes of kernels which pair with low-discrepancy sequences to accelerate kernel computations. This includes new higher-order digitally shift invariant kernels. The accelerated methods utilize modified fast Fourier transforms and the fast Walsh-Hadamard transform. The presented material has been implemented into the accessible open-source Python library QMCPy.
Acknowledgements.
@@
Thank you to Fred J Hickernell, Sou-Cheng T Choi, and Aadit Jain for helpful comments and prototyping. The author would like to acknowledge the support of NSF Grant DMS-2316011.
Appendix A Proofs of theorems

### v2

@@ 6.3 Fast kernel computations and efficient eigenvalue updates
    where ${\mathsf{V}}_{m}\overline{{\mathsf{V}}_{m}}={\mathsf{I}}$ and the first column of ${\mathsf{V}}_{m}$, denoted ${\boldsymbol{v}}_{m1}$, is the constant $1/\sqrt{n}$.
  2. 2.
    ${\mathsf{V}}_{m}\in{\mathbb{C}}^{n\times n}$ satisfies ${\mathsf{V}}_{0}=1$ and
@@
    ${\mathsf{V}}_{m+1}=\begin{pmatrix}{\mathsf{V}}_{m}&{\mathsf{V}}_{m}\\ {\mathsf{V}}_{m}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})}&-{\mathsf{V}}_{m}\overline{\mathrm{diag}({\widetilde{\boldsymbol{w}}}_{m})}\end{pmatrix}/\sqrt{2}.$ (8)
  3. 3.
    ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ are each computable at cost ${\mathcal{O}}(2^{m}m)={\mathcal{O}}(n\log n)$.
@@
For the case of lattice points with SI kernels, ${\mathsf{V}}_{m}={\mathsf{T}}_{m}/\sqrt{2^{m}}$ is the scaled and permuted Fourier matrix and ${\widetilde{\boldsymbol{w}}}_{m}=\{\exp(-\pi\sqrt{-1}/2^{m}i)\}_{i=0}^{2^{m}-1}$, so ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ can be computed using the FFTBR and IFFTBR respectively. For the case of digital nets with DSI kernels, ${\mathsf{V}}_{m}={\mathsf{H}}_{m}/\sqrt{n}$ is the real symmetric Hadamard matrix and ${\widetilde{\boldsymbol{w}}}_{m}=\{1\}_{i=0}^{2^{m}-1}$, so both ${\mathsf{V}}_{m}{\boldsymbol{y}}$ and $\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$ can be computed using a FWHT.
With ${\mathsf{\Lambda}}_{m}=\mathrm{diag}({\boldsymbol{\lambda}}_{,})$, we have
${\boldsymbol{\lambda}}_{m}=\sqrt{n}{\mathsf{\Lambda}}_{m}\overline{{\boldsymbol{v}}_{m1}}=\sqrt{n}\overline{{\mathsf{V}}_{m}}\left({\mathsf{V}}_{m}{\mathsf{\Lambda}}_{m}\overline{{\boldsymbol{v}}_{m1}}\right)=\sqrt{n}\overline{{\mathsf{V}}_{m}}{\boldsymbol{k}}_{m1}$
which can be computed at ${\mathcal{O}}(n\log n)$ cost and only requires storing the first column of ${\mathsf{K}}^{(m)}$ which we denote by ${\boldsymbol{k}}_{m1}$. Moreover,
${\mathsf{K}}{\boldsymbol{y}}={\mathsf{V}}_{m}({\widetilde{\boldsymbol{y}}}\odot{\boldsymbol{\lambda}}_{m})\qquad\text{and}\qquad{\mathsf{K}}^{-1}{\boldsymbol{y}}={\mathsf{V}}_{m}({\widetilde{\boldsymbol{y}}}./{\boldsymbol{\lambda}}_{m})$
@@
may each be evaluated at cost ${\mathcal{O}}(n\log n)$ where $\odot$ denotes the Hadamard (component-wise) product and ././ denotes component-wise division.
The following code implements these fast Gram matrix operations for the SI-lattice variant. It also includes an application of (8) to update transformed values, i.e., to compute ${\widetilde{\boldsymbol{y}}}_{\mathrm{full}}=\overline{{\mathsf{V}}_{m+1}}\begin{pmatrix}{\boldsymbol{y}}\\ {\boldsymbol{y}}_{\mathrm{new}}\end{pmatrix}$ given ${\widetilde{\boldsymbol{y}}}=\overline{{\mathsf{V}}_{m}}{\boldsymbol{y}}$. An analogous code for the DSI-digital net variant is given in Listing LABEL:code:dsi_dnet.
Listing 5: Efficient Gram matrix operations and transform updates for the SI-lattice variant.
d = 3 # dimension
m = 10 # will generate 2^m points
n = 2**m # number of points
lattice = qp.Lattice(d) # defaults to radical inverse order
kernel = qp.KernelShiftInvar(
d, # dimension
alpha = [1,2,3], # per-dimension smoothness parameters
lengthscales = [1, 1/2, 1/4]) # per-dimension product weights
x = lattice(n_min=0,n_max=n) # shape=(n, d) lattice points
y = np.random.rand(n) # shape=(n,) random uniforms
# fast matrix multiplication and linear system solve
k1 = kernel(x,x[0]) # shape=(n,) first column of Gram matrix
lam = np.sqrt(n)*qp.fftbr(k1) # vector of eigenvalues
yt = qp.fftbr(y)
u = qp.ifftbr(yt*lam) # fast matrix multiplication
v = qp.ifftbr(yt/lam) # fast linear system solve
# efficient fast transform updates
ynew = np.random.rand(n) # shape=(n,) new random uniforms
omega = qp.omega_fftbr(m) # shape=(n,)
ytnew = qp.fftbr(ynew) # shape=(n,)
ytfull1 = yt+omega*ytnew # shape=(n,) first half of ytfull
y2full2 = yt-omega*ytnew # shape=(n,) second half of ytfull
ytfull = 1/np.sqrt(2)*np.hstack([ytfull1,y2full2]) # shape=(2n,)
7 Numerical experiments
@@
All numerical experiments were carried out on a 2023 MacBook Pro with an M3 Max processor. Figure 4 compares the wall-clock time required to generate point sets and perform fast transforms for high dimensions and/or number of randomizations. Following the theory, our implementation scales linearly in the number of dimensions and randomizations. For large numbers of randomizations, our vectorized QMCPy generators significantly outperform the looped implementations in SciPy and PyTorch. For fast transforms, our FFTBR and IFFTBR implementations are the same speed as SciPy’s FFT and IFFT algorithms. Our FWHT implementation is significantly faster than the SymPy version, especially when applying the FWHT to multiple sequences simultaneously. IID points, randomly shifted lattices, and digital nets with LMS and digital shifts (including higher-order versions) are the fastest sequences to generate. Digital nets in base $b=2$ exploit Gray code order, integer storage of bit-vectors, and exclusive or (XOR) operations to perform digital addition. Halton point sets are slower to generate as they cannot exploit these advantages. NUS, especially higher-order versions, are significantly slower to generate than LMS randomizations. Even so, higher-order LMS scrambling with digital shifts are sufficient to achieve higher-order RMSE convergence as we show in the next experiment.
Figure 4: Comparison of time required to generate point sets and perform fast transforms. For IID and randomized LD generators, we vary the number of randomizations $R$, the number of dimensions $d$, and the sequence size $n$. Timings include initialization, and a new point set is generating for each plotted $(R,n,d)$. Fast transforms are applied to $r$ sequences of size $n$ in a vectorized fashion.
Figure 5 plots the RMSE convergence of RQMC methods applied to the integrands described below.
Simple function, $d=1$,
has $f(x)=xe^{x}-1$. This was used in [Dick, 2011] where higher-order digital net scrambling was first proposed.
Simple function, $d=2$,
has $f({\boldsymbol{x}})=x_{2}e^{x_{1}x_{2}}/(e-2)-1$. This was also considered in [Dick, 2011].
Oakley & O’Hagan, $d=2$,
has $f({\boldsymbol{x}})=g(({\boldsymbol{x}}-1/2)/50)$ for $g({\boldsymbol{t}})=5+t_{1}+t_{2}+2\cos(t_{1})+2\cos(t_{2})$, see [Oakley and O’hagan, 2002] or the VLSE [^17] https://www.sfu.ca/~ssurjano/uq.html (Virtual Library of Simulation Experiments).
G-Function, $d=3$,
has $f({\boldsymbol{x}})=\prod_{j=1}^{d}\frac{\lvert 4x_{j}-2\rvert-a_{j}}{1+a_{j}}$ with $a_{j}=(j-2)/2$ for $1\leq j\leq d$, see [Crestaux et al., 2007, Marrel et al., 2009] or the VLSE.
Oscillatory Genz, $d=3$,
has $f({\boldsymbol{x}})=\cos\left(-\sum_{j=1}^{d}c_{j}x_{j}\right)$ with coefficients of the third kind $c_{j}=4.5{\widetilde{c}}_{j}/\sum_{j=1}^{d}{\widetilde{c}}_{j}$ where ${\widetilde{c}}_{j}=\exp\left(j\log\left(10^{-8}\right)/d\right)$. This is a common test function for uncertainty quantification which is available in the Dakota [^18] https://dakota.sandia.gov/ software [Adams et al., 2020] among others.
Corner-peak Genz, $d=3$,
has $f({\boldsymbol{x}})=\left(1+\sum_{j=1}^{d}c_{j}x_{j}\right)^{-(d+1)}$ with coefficients of the second kind $c_{j}=0.25{\widetilde{c}}_{j}/\sum_{j=1}^{d}{\widetilde{c}}_{j}$ where ${\widetilde{c}}_{j}=1/j^{2}$. This is also available in Dakota.
@@
Figure 5: The RMSE of the RQMC estimate for a few different integrands. Higher-order digital nets with linear matrix scrambling and digital shifts achieve higher-order convergence and significantly outperform both IID points and randomly shifted lattices.
In the following code, we use QMCPy to build an RQMC estimate to the mean of the Corner-peak Genz function in $d=50$ dimensions. Using $R$ replications and a fixed number of points $n$, we construct a Student’s-$t$ confidence interval as in (2) and the discussion thereafter. SciPy is used to compute the necessary quantile. Listing LABEL:code:genz_ex_2 gives an example of an adaptive version of this algorithm which automatically selects the number of points $n$ required to meet a user-specified absolute error tolerance.
Listing 6: Randomized QMC approximation and error estimation for a fixed number of points.
import scipy.stats
def gen_corner_peak_2(x):
d = x.shape[-1] # x.shape=(...,n,d), e.g., (n,d) or (R,n,d)
c_tilde = 1/np.arange(1,d+1)**2
c = 0.25*c_tilde/np.sum(c_tilde)
y = (1+np.sum(c*x,axis=-1))**(-(d+1))
return y # y.shape=(...,n), e.g., (n,) or (R,n)
R = 10 # number of randomizations
n = 2**15 # number of points
d = 50 # dimension
dnb2 = qp.DigitalNet(dimension=d, replications=R, seed=7, alpha=3)
x = dnb2(n) # x.shape=(R,n,d)
y = gen_corner_peak_2(x) # y.shape=(R,n)
muhats_r = np.mean(y,axis=1) # muhats_r.shape=(R,)
muhat = np.mean(muhats_r) # muhat is a scalar
print(muhat)
""" 0.014936813948394042 """
alpha = 0.01 # uncertainty level
t_star = -scipy.stats.t.ppf(alpha/2,df=R-1) # quantile
sigmahat = np.std(muhats_r,ddof=1) # unbiased estimate
std_error = t_star*sigmahat/np.sqrt(R)
print(std_error)
""" 5.247445301861484e-07 """
conf_int = [muhat-std_error,muhat+std_error]
8 Conclusion
@@
This work has reviewed our QMCPy implementations of routines for generating randomized LD sequences and applying them to fast kernel methods. We presented lattices, digital nets, and Halton point sets with randomizations spanning shifts, digital permutations, digital shifts, LMS, and NUS. Higher-order scramblings for digital nets were also considered, including higher-order LMS and NUS. We described classes of kernels which pair with LD sequences to enable fast kernel computations. These included known SI kernels as well as new DSI kernels of higher-order smoothness. The fast kernel methods utilized the fast transform algorithms FFTBR, IFFTBR, and FWHT. We showed our implementations achieves state-of-the-art speed, accuracy, and flexibility.
Acknowledgements
@@
Thank you to Fred J Hickernell, Sou-Cheng T Choi, and Aadit Jain for helpful comments and prototyping. The author would like to acknowledge the support of NSF Grant DMS-2316011 and the DOE Office of Science Graduate Student Research (SCGSR) Program.
Appendix A Extra Code Snippets
Listing 7: Efficient Gram matrix operations and transform updates for the DSI-digital net variant.
d = 3 # dimension
m = 10 # will generate 2^m points
n = 2**m # number of points
dnb2 = qp.DigitalNetB2(d) # defaults to radical inverse order
kernel = qp.KernelDigShiftInvar(
d, # dimension
t = dnb2.t, # number of bits in integer representation of points
alpha = [1,2,3], # per-dimension smoothness parameters
lengthscales = [1, 1/2, 1/4]) # per-dimension product weights
x = dnb2(n_min=0,n_max=n) # shape=(n, d) digital net
y = np.random.rand(n) # shape=(n,) random uniforms
# fast matrix multiplication and linear system solve
k1 = kernel(x,x[0]) # shape=(n,) first column of Gram matrix
lam = np.sqrt(n)*qp.fwht(k1) # vector of eigenvalues
yt = qp.fwht(y)
u = qp.fwht(yt*lam) # fast matrix multiplication
v = qp.fwht(yt/lam) # fast linear system solve
# efficient fast transform updates
ynew = np.random.rand(n) # shape=(n,) new random uniforms
omega = qp.omega_fwht(m) # shape=(n,)
ytnew = qp.fwht(ynew) # shape=(n,)
ytfull1 = yt+omega*ytnew # shape=(n,) first half of ytfull
y2full2 = yt-omega*ytnew # shape=(n,) second half of ytfull
ytfull = 1/np.sqrt(2)*np.hstack([ytfull1,y2full2]) # shape=(2n,)
Listing 8: Adaptive randomized QMC approximation and error estimation.
def gen_corner_peak_2(x):
d = x.shape[-1] # x.shape=(...,n,d), e.g., (n,d) or (R,n,d)
c_tilde = 1/np.arange(1,d+1)**2
c = 0.25*c_tilde/np.sum(c_tilde)
y = (1+np.sum(c*x,axis=-1))**(-(d+1))
return y # y.shape=(...,n), e.g., (n,) or (R,n)
R = 10
d = 50
dnb2 = qp.DigitalNet(dimension=d, replications=R, seed=7, alpha=3)
true_measure = qp.Uniform(dnb2, lower_bound=0, upper_bound=1)
integrand = qp.CustomFun(true_measure=true_measure, g=gen_corner_peak_2)
# equivalent to
# integrand = qp.Genz(dnb2, kind_func="CORNER PEAK", kind_coeff=2)
qmc_algo = qp.CubQMCRepStudentT(integrand, abs_tol=1e-4)
solution,data = qmc_algo.integrate() # run adaptive QMC algorithm
print(solution)
""" 0.014950908095474802 """
conf_int = [data.comb_bound_low,data.comb_bound_high]
std_error = (conf_int[1]-conf_int[0])/2
print(std_error)
""" 2.7968149935497788e-05 """
print(data)
Data (Data)
solution 0.015
comb_bound_low 0.015
comb_bound_high 0.015
comb_bound_diff 5.59e-05
n 10240
n_rep 2^(10)
time_integrate 0.019
CubQMCRepStudentT (AbstractStoppingCriterion)
alpha 0.010
abs_tol 1.00e-04
rel_tol 0
n_init 2^(8)
DigitalNetB2 (AbstractLDDiscreteDistribution)
d 50
replications 10
randomize LMS DS
gen_mats_source joe_kuo.6.21201.txt
order RADICAL INVERSE
t 63
alpha 3
entropy 7
Appendix B Proofs of theorems